OpenStudy (anonymous):

lim x goes to (pi/2), [ln(sinx) + x - (pi/2)] / [cosx]

OpenStudy (anonymous):

\[\lim_{x\rightarrow \frac{\pi}{2}} \frac{\ln(\sin(x))+x-\frac{\pi}{2}}{\cos(\frac{\pi}{2})}\] this one?

OpenStudy (anonymous):

or rather i meant \[\lim_{x\rightarrow \frac{\pi}{2}} \frac{\ln(\sin(x))+x-\frac{\pi}{2}}{\cos(x)}\]

OpenStudy (anonymous):

yes thats the one, i got -2 as an answer is that right? i used l'hospitals rule

OpenStudy (anonymous):

get 0/0 straight up l'hopital problem. take the derivative top and bottom get \[\frac{\frac{\cos(x)}{\sin(x)}+1}{-\sin(x)}\]

OpenStudy (anonymous):

replace x by \[\frac{\pi}{2}\] get -1

OpenStudy (anonymous):

not sure where the -2 came from unless you forgot the chain rule when taking the derivative of \[\ln(\sin(x))\]